# FALCON: Learning Force-Adaptive Humanoid Loco-Manipulation

### v1

@@ Abstract
Humanoid loco-manipulation holds transformative potential for daily service and industrial tasks, yet achieving precise, robust whole-body control with 3D end-effector force interaction remains a major challenge. Prior approaches are often limited to lightweight tasks or quadrupedal/wheeled platforms. To overcome these limitations, we propose FALCON, a dual-agent reinforcement-learning-based framework for robust force-adaptive humanoid loco-manipulation. FALCON decomposes whole-body control into two specialized agents: (1) a lower-body agent ensuring stable locomotion under external force disturbances, and (2) an upper-body agent precisely tracking end-effector positions with implicit adaptive force compensation. These two agents are jointly trained in simulation with a force curriculum that progressively escalates the magnitude of external force exerted on the end effector while respecting torque limits. Experiments demonstrate that, compared to the baselines, FALCON achieves 2$\times$ more accurate upper-body joint tracking, while maintaining robust locomotion under force disturbances and achieving faster training convergence. Moreover, FALCON enables policy training without embodiment-specific reward or curriculum tuning. Using the same training setup, we obtain policies that are deployed across multiple humanoids, enabling forceful loco-manipulation tasks such as transporting payloads (0-20N force), cart-pulling (0-100N), and door-opening (0-40N) in the real world.
@@ 1 Introduction
Humanoid robots have demonstrated remarkable progress in locomotion and manipulation [1, 2, 3, 4, 5, 6]. However, extending these capabilities to forceful loco-manipulation remains fundamentally challenging. Tasks such as door opening, highlighted in the 2015 DARPA Challenge [7], require not only precise manipulation under dynamic, multi-directional forces but also maintaining lower-body stability throughout the interaction. Meeting these demands calls for humanoid systems that can flexibly adapt to varying payloads and contact forces without compromising overall precision and robustness in loco-manipulation.
Reinforcement Learning (RL) has achieved impressive results for humanoid whole-body control [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27], yet existing RL approaches succeed mostly on lightweight tasks but do not consider significant interaction force during loco-manipulation tasks. Currently, there are two main paradigms: (1) Lower-RL-Upper-IK, which applies RL to lower-body locomotion while using kinematic solvers for upper-body control [8, 9], lacks whole-body dynamics modeling for forceful interaction and has limited whole-body coordination; (2) Monolithic-Whole-body-RL, which directly learns to control all degrees of freedom [10, 11], suffers from inefficient exploration as a single policy must simultaneously learn weakly correlated locomotion and manipulation skills. Although some advances have been made in force adaptation for quadrupeds [12, 13, 14, 28, 29], humanoids pose extra challenges like instability, higher complexity, and stricter torque limits, especially in certain joint configurations.
@@ 2.1 Humanoid Loco-Manipulation
Humanoid loco-manipulation remains a challenging control problem in robotics. While traditional model-based methods (e.g., simplified dynamics models and MPC) [20, 13, 21, 22, 23, 24, 25, 30] offer real-time planning, their reliance on manual design limits flexibility and generalizability. In contrast, learning-based methods—particularly sim-to-real RL—have demonstrated promising results in versatile loco-manipulation tasks [31, 17, 11, 8, 9, 32, 33]. For humanoids, two primary paradigms have emerged: Lower-RL-Upper-IK and Monolithic-Whole-body-RL. For Lower-RL-Upper-IK, Lu et al. [8] introduce PMP, which uses inverse kinematics (IK) and PD control for upper body control while locomotion is trained and conditioned on a Conditional Variational Autoencoder (CVAE) representing upper-body motions. Then, Ben et al. [9] propose HOMIE that follows the same decoupling framework but introduces an exoskeleton-based cockpit for more intuitive human teleoperation. For Monolithic-Whole-body-RL, Dao et al. [32] adopted a unified RL approach for box pick-and-place tasks, training distinct skills (e.g., lifting, walking, stance) and orchestrating them via a finite state machine. He et al. [16, 17] and Ji et al. [10] employ a teacher-student training framework to mimic human motions for loco-manipulation tasks.
Despite these advances, few RL methods address significant unknown force disturbances on the EEs for humanoid loco-manipulation, and both paradigms exhibit critical shortcomings accordingly. Lower-RL-Upper-IK approaches suffer from delayed force compensation for upper-body control. Monolithic-Whole-body-RL methods face sample inefficiency from coarsely related task objectives between upper-body manipulation and lower-body locomotion, often leading to overfitting and the behavioral dominance of either upper or lower body. In this study, inspired by [34, 35, 36], we propose FALCON, a dual-agent RL framework employing task-specific reward formulations for upper-lower body decomposition. Unlike separately trained architectures, the two agents in FALCON are jointly trained with shared proprioception and commands, allowing mutual awareness of each other’s behaviors. This joint training prevents the agents from adapting in isolation and enables coordinated responses to external forces that affect the full-body dynamics.
@@ 2.2 Forceful Interaction in Legged Robots
Forceful interaction has been extensively studied for quadrupedal robots with mounted arms, through model-based approaches—particularly MPC combined with force planning and control for robust and adaptive locomotion and manipulation [12, 13, 37]. Recent advances in RL have further enhanced adaptability, enabling quadrupeds to learn adaptive and agile force interactions including impedance control [38] and aggressive force adaptation [14]. For humanoids, forceful interaction presents significantly greater challenges due to their more complex dynamics and stringent joint limits. Unlike quadrupeds with centralized mass distributions, humanoids exhibit coupled dynamics between their upper and lower bodies, making force adaptation particularly difficult. Recent model-based approaches have demonstrated force control for heavy-duty tasks [39, 22], but these require prior knowledge of manipulated objects’ mass, center of mass (CoM), or pre-defined force trajectories, limiting their applicability to unknown disturbances. While some works have attempted explicit force estimation for humanoids [40], they are restricted to quasi-static scenarios and cannot handle force adaptation in dynamic loco-manipulation scenarios.
@@ 3 FALCON: Force-Adaptive Humanoid Loco-Manipulation
Figure 2: Overview of FALCON. (a) Two agents with different sub-tasks are jointly trained with shared whole-body proprioception. During training, we apply 3D external forces bounded by upper-body joint torque limits on the end-effectors; (b) FALCON is deployed with either teleoperation or an autonomy pipeline including FoundationPose [41] for pose estimation and motion planning.
@@ 3.1 Dual-Agent Learning Framework
[…]
These two policy parameters $\theta_{l}$ and $\theta_{u}$ are updated via proximal policy optimization (PPO [42]):
[…]
where $\gamma$ is the discount factor. The upper-body target joint angles $\mathbf{q}^{\text{upper}}_{t}$ (target joints of shoulders, elbows, wrists) are randomly sampled from the AMASS dataset [43] during training, and calculated via IK during deployment. The combined action from the two agents $\boldsymbol{a}_{t}=[\boldsymbol{a}^{l}_{t};\boldsymbol{a}^{u}_{t}]$ is sent to a joint-level PD controller. As the real-world humanoid control is inherently partially observable, we adopt asymmetric actor-critic training, where critics additionally access privileged information including root linear velocities and EE forces $\boldsymbol{F}^{ee}_{t}$ during training but not during deployment. Detailed reward designs and domain randomization during training are provided in Section A.1 and Section A.2.
@@ 3.2 Torque-Limit-Aware 3D Force Curriculum
Before applying forces, we first need to estimate the maximum forces that we can exert on the left or right end-effector. Given the left or right end-effector Jacobian $\boldsymbol{J}_{EE}\in\mathbb{R}^{3\times\frac{n_{u}}{2}}$ at its Center of Mass (CoM), their joint torque limit $\boldsymbol{\tau}^{\lim}\in\mathbb{R}^{3\times\frac{n_{u}}{2}}$ (with $\boldsymbol{\tau}^{\lim}\geq\boldsymbol{0}$), and the gravity compensation torque $\boldsymbol{\tau}^{g}\in\mathbb{R}^{3\times\frac{n_{u}}{2}}$ which satisfies $-\boldsymbol{\tau}^{\lim}\leq\boldsymbol{\tau}^{g}\leq\boldsymbol{\tau}^{\lim}$ to ensure feasibility, we estimate the maximum and minimum admissible forces $\boldsymbol{f}^{\max},\boldsymbol{f}^{\min}$ along each Cartesian axis $i\in{x,y,z}$ by analyzing the worst-case joint torque induced by a unit force applied in each direction. The element-wise force bound can be computed in parallel as:
$-\boldsymbol{\tau}^{\lim}\leq\boldsymbol{\tau}^{g}+\boldsymbol{J}_{EE}^{T}\boldsymbol{f}^{ee}\leq\boldsymbol{\tau}^{\lim}$ (3)
@@
where $J_{EE}^{ji}$ denotes the $(j,i)$-th element of the end-effector Jacobian matrix, and $\epsilon$ is a small positive constant to prevent division by zero. After that, we sample the relatively ratio $\boldsymbol{\gamma}=[\gamma_{x},\gamma_{y},\gamma_{z}]$ among x, y and z axis through Dirichlet Distribution [44], which satisfy $\sum_{i\in\{x,y,z\}}\gamma_{i}=1$. The feasible applied force will be uniformly sampled within the estimated range and expressed as:
@@
Learning-based force adaptation leverages proprioceptive history to implicitly compensate for external forces, removing the need for explicit force estimation [40] or sensing [45]. To improve robustness to variations in end-effector (EE) contact points—which alter the torque mapping via the EE Jacobian—we randomize force application along the EE link, from the wrist yaw to the distal segment, as illustrated in Figure 2.
@@ 4.1 Evaluation Criterion
To evaluate the performance of the learned low-body locomotion and upper-body manipulation capabilities, we consider the following metrics under dynamically unknown and 3D EE forces $\textbf{F}_{t}\in\mathbb{R}^{3}$, given a sequence of target upper-body joints $\boldsymbol{q}^{\text{upper*}}_{t}$, target root velocities $\boldsymbol{v}^{\text{lin,ang*}}_{t}$ and stance signal $\phi_{t}^{\text{stance}}$, where $t=1,2,...,T$ and $T$ is the sequence length:
(ii) Upper-Body Joints Tracking Error: $*E_{\text{tracking}}^{\text{upper}}(\boldsymbol{q}^{\text{upper*}}_{t})=\frac{1}{T}\sum_{t=1}^{T}\left|\boldsymbol{q}_{t}^{\text{upper}}-\boldsymbol{q}_{t}^{\text{upper*}}\right|$
(iii) Root Velocity Tracking Error: $*E_{\text{tracking}}^{\text{root}}(\boldsymbol{v}_{t}^{\text{lin,ang*}})=\frac{1}{T}\sum_{t=1}^{T}\left|\boldsymbol{v}_{t}^{\text{lin,ang}}-\boldsymbol{v}_{t}^{\text{lin,ang*}}\right|$
@@ 4.2 Baselines
    Upper-PD-w/o-Force-Curr.: A baseline following [8, 9], using PD control for upper-body joint tracking without force randomization.
@@
    Upper-PD-ID: Extends (a) with a learned force estimator [38] and inverse-dynamics-based torque compensation under quasi-static assumptions (details in Section A.3).
@@
    Monolithic-WB-RL-w/o-Force-Curr.: Built upon prior designs [11, 15], a single agent is trained with the same goal commands as FALCON, but without applying any force during training.
@@ 4.3 Simulation Results
We evaluate FALCON and baselines on 252 ACCAD [46] motion targets under three force levels: (i) No-Force ($\alpha_{g}=0$), (ii) Middle-Force ($\alpha_{g}=0.5$), and (iii) Large-Force ($\alpha_{g}=1.0$), applied to both end-effectors. As shown in Table 1, across all settings, FALCON with force curriculum achieves the lowest tracking errors in both upper-body motion ($E^{\text{upper}}_{\text{tracking}}$) and root velocity ($E^{\text{root}}_{\text{tracking}}$), demonstrating robust manipulation under disturbance. Under L-Force, it reduces upper-body error to 0.37, outperforming PID-Force-Curr. (0.60) and M-WB-RL (0.73). Root error remains low at 0.45, indicating stable locomotion. While force curriculum benefits all methods, FALCON gains most due to its decomposed learning structure.
@@ 4.4 Real-World Quantatitive Tracking Results
We evaluate FALCON on Unitree G1 with each hand loaded with 1.2kg payload in a real-world task, which is walking at $(0.5,0.0)$m/s with zero angular velocity, fixed height and waist, and keeping the upper body in its default position. We compare against two baselines: (i) Upper-PD with Force Curriculum, and (ii) Monolithic-WB-RL with Force Curriculum. As shown in Table 4, FALCON achieves the lowest tracking errors, and perform stable and natural motion in heavy-duty loco-manipulation.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Humanoid robots have demonstrated remarkable progress in locomotion and manipulation [gu2025humanoid, darvish2023teleoperation, atlas, unitreeg1, figureai2024, Booster2025T1]. However, extending these capabilities to forceful loco-manipulation remains fundamentally challenging. Tasks such as door opening, highlighted in the 2015 DARPA Challenge [dappa], require not only precise manipulation under dynamic, multi-directional forces but also maintaining lower-body stability throughout the interaction. Meeting these demands calls for humanoid systems that can flexibly adapt to varying payloads and contact forces without compromising overall precision and robustness in loco-manipulation.
Reinforcement Learning (RL) has achieved impressive results for humanoid whole-body control [lu2024mobile, homie, ji2024exbody2, cheng2024expressive, sombolestan2023hierarchical, sombolestan2024adaptive, fey2025bridging, he2024hover, he2024learning, he2024omnih2o, he2025asap, zhuang2025embrace, dantec2021whole, li2023dynamic, murooka2021humanoid, bouyarmane2018quadratic, di2018dynamic, kajita2003biped, fu2024humanplus, zhang2024wococo], yet existing RL approaches succeed mostly on lightweight tasks but do not consider significant interaction force during loco-manipulation tasks. Currently, there are two main paradigms: (1) Lower-RL-Upper-IK, which applies RL to lower-body locomotion while using kinematic solvers for upper-body control [lu2024mobile, homie], lacks whole-body dynamics modeling for forceful interaction and has limited whole-body coordination; (2) Monolithic-Whole-body-RL, which directly learns to control all degrees of freedom [ji2024exbody2, cheng2024expressive], suffers from inefficient exploration as a single policy must simultaneously learn weakly correlated locomotion and manipulation skills. Although some advances have been made in force adaptation for quadrupeds [sombolestan2023hierarchical, sombolestan2024adaptive, fey2025bridging, cheng2025rambo, zhong2025bridging], humanoids pose extra challenges like instability, higher complexity, and stricter torque limits, especially in certain joint configurations.
@@ 2.1 Humanoid Loco-Manipulation
Humanoid loco-manipulation remains a challenging control problem in robotics. While traditional model-based methods (e.g., simplified dynamics models and MPC) [dantec2021whole, sombolestan2024adaptive, li2023dynamic, murooka2021humanoid, bouyarmane2018quadratic, di2018dynamic, kajita2003biped, xue2024full] offer real-time planning, their reliance on manual design limits flexibility and generalizability. In contrast, learning-based methods—particularly sim-to-real RL—have demonstrated promising results in versatile loco-manipulation tasks [zhang2024catchit, he2024omnih2o, cheng2024expressive, lu2024mobile, homie, dao2024sim, liu2024opt2skill]. For humanoids, two primary paradigms have emerged: Lower-RL-Upper-IK and Monolithic-Whole-body-RL. For Lower-RL-Upper-IK, lu2024mobile introduce PMP, which uses inverse kinematics (IK) and PD control for upper body control while locomotion is trained and conditioned on a Conditional Variational Autoencoder (CVAE) representing upper-body motions. Then, homie propose HOMIE that follows the same decoupling framework but introduces an exoskeleton-based cockpit for more intuitive human teleoperation. For Monolithic-Whole-body-RL, dao2024sim adopted a unified RL approach for box pick-and-place tasks, training distinct skills (e.g., lifting, walking, stance) and orchestrating them via a finite state machine. he2024learning, he2024omnih2o and ji2024exbody2 employ a teacher-student training framework to mimic human motions for loco-manipulation tasks.
Despite these advances, few RL methods address significant unknown force disturbances on the EEs for humanoid loco-manipulation, and both paradigms exhibit critical shortcomings accordingly. Lower-RL-Upper-IK approaches suffer from delayed force compensation for upper-body control. Monolithic-Whole-body-RL methods face sample inefficiency from coarsely related task objectives between upper-body manipulation and lower-body locomotion, often leading to overfitting and the behavioral dominance of either upper or lower body. In this study, inspired by [gronauer2022multi, zhang2021multi, wang2024learning], we propose FALCON, a dual-agent RL framework employing task-specific reward formulations for upper-lower body decomposition. Unlike separately trained architectures, the two agents in FALCON are jointly trained with shared proprioception and commands, allowing mutual awareness of each other’s behaviors. This joint training prevents the agents from adapting in isolation and enables coordinated responses to external forces that affect the full-body dynamics.
@@ 2.2 Forceful Interaction in Legged Robots
Forceful interaction has been extensively studied for quadrupedal robots with mounted arms, through model-based approaches—particularly MPC combined with force planning and control for robust and adaptive locomotion and manipulation [sombolestan2023hierarchical, sombolestan2024adaptive, rigo2024hierarchical]. Recent advances in RL have further enhanced adaptability, enabling quadrupeds to learn adaptive and agile force interactions including impedance control [portela2024learning] and aggressive force adaptation [fey2025bridging]. For humanoids, forceful interaction presents significantly greater challenges due to their more complex dynamics and stringent joint limits. Unlike quadrupeds with centralized mass distributions, humanoids exhibit coupled dynamics between their upper and lower bodies, making force adaptation particularly difficult. Recent model-based approaches have demonstrated force control for heavy-duty tasks [li2023kinodynamics, murooka2021humanoid], but these require prior knowledge of manipulated objects’ mass, center of mass (CoM), or pre-defined force trajectories, limiting their applicability to unknown disturbances. While some works have attempted explicit force estimation for humanoids [mattioli2016interaction], they are restricted to quasi-static scenarios and cannot handle force adaptation in dynamic loco-manipulation scenarios.
@@ 3 FALCON: Force-Adaptive Humanoid Loco-Manipulation
Figure 2: Overview of FALCON. (a) Two agents with different sub-tasks are jointly trained with shared whole-body proprioception. During training, we apply 3D external forces bounded by upper-body joint torque limits on the end-effectors; (b) FALCON is deployed with either teleoperation or an autonomy pipeline including FoundationPose [wen2024foundationpose] for pose estimation and motion planning.
@@ 3.1 Dual-Agent Learning Framework
$r^{l}_{t}=\mathcal{R}^{l}(\boldsymbol{s}^{p}_{t},\mathcal{G}^{l}_{t})~\text{(locomotion)}\quad r^{u}_{t}=\mathcal{R}^{u}(\boldsymbol{s}^{p}_{t},\mathcal{G}^{u}_{t})~\text{(manipulation)}$ (1)
These two policy parameters $\theta_{l}$ and $\theta_{u}$ are updated via proximal policy optimization (PPO [PPO]):
$\max_{\theta_{l}}\mathbb{E}\left[\sum_{t=1}^{T}\gamma^{t-1}r^{l}_{t}\right]~\text{(Lower-body)}\quad\max_{\theta_{u}}\mathbb{E}\left[\sum_{t=1}^{T}\gamma^{t-1}r^{u}_{t}\right]~\text{(Upper-body)}$ (2)
where $\gamma$ is the discount factor. The upper-body target joint angles $\mathbf{q}^{\text{upper}}_{t}$ (target joints of shoulders, elbows, wrists) are randomly sampled from the AMASS dataset [AMASS] during training, and calculated via IK during deployment. The combined action from the two agents $\boldsymbol{a}_{t}=[\boldsymbol{a}^{l}_{t};\boldsymbol{a}^{u}_{t}]$ is sent to a joint-level PD controller. As the real-world humanoid control is inherently partially observable, we adopt asymmetric actor-critic training, where critics additionally access privileged information including root linear velocities and EE forces $\boldsymbol{F}^{ee}_{t}$ during training but not during deployment. Detailed reward designs and domain randomization during training are provided in Section A.1 and Section A.2.
@@ 3.2 Torque-Limit-Aware 3D Force Curriculum
Before applying forces, we first need to estimate the maximum forces that we can exert on the left or right end-effector. Given the left or right end-effector Jacobian $\boldsymbol{J}_{EE}\in\mathbb{R}^{3\times\frac{n_{u}}{2}}$ at its Center of Mass (CoM), their joint torque limit $\boldsymbol{\tau}^{\lim}\in\mathbb{R}^{3\times\frac{n_{u}}{2}}$ (with $\boldsymbol{\tau}^{\lim}\geq\boldsymbol{0}$), and the gravity compensation torque $\boldsymbol{\tau}^{g}\in\mathbb{R}^{3\times\frac{n_{u}}{2}}$ which satisfies $\boldsymbol{\tau}^{\lim}\leq\boldsymbol{\tau}^{g}\leq\boldsymbol{\tau}^{\lim}$ to ensure feasibility, we estimate the maximum and minimum admissible forces $\boldsymbol{f}^{\max},\boldsymbol{f}^{\min}$ along each Cartesian axis $i\in{x,y,z}$ by analyzing the worst-case joint torque induced by a unit force applied in each direction. The element-wise force bound can be computed in parallel as:
$\boldsymbol{\tau}^{\lim}\leq\boldsymbol{\tau}^{g}+\boldsymbol{J}_{EE}^{T}\boldsymbol{f}^{ee}\leq\boldsymbol{\tau}^{\lim}$ (3)
@@
where $J_{EE}^{ji}$ denotes the $(j,i)$-th element of the end-effector Jacobian matrix, and $\epsilon$ is a small positive constant to prevent division by zero. After that, we sample the relatively ratio $\boldsymbol{\gamma}=[\gamma_{x},\gamma_{y},\gamma_{z}]$ among x, y and z axis through Dirichlet Distribution [ng2011dirichlet], which satisfy $\sum_{i\in\{x,y,z\}}\gamma_{i}=1$. The feasible applied force will be uniformly sampled within the estimated range and expressed as:
@@
Learning-based force adaptation leverages proprioceptive history to implicitly compensate for external forces, removing the need for explicit force estimation [mattioli2016interaction] or sensing [guo2024flying]. To improve robustness to variations in end-effector (EE) contact points—which alter the torque mapping via the EE Jacobian—we randomize force application along the EE link, from the wrist yaw to the distal segment, as illustrated in Figure 2.
@@ 4.1 Evaluation Criterion
To evaluate the performance of the learned low-body locomotion and upper-body manipulation capabilities, we consider the following metrics under dynamically unknown and 3D EE forces $\textbf{F}_{t}\in\mathbb{R}^{3}$, given a sequence of target upper-body joints $\boldsymbol{q}^{\text{upper*}}_{t}$, target root velocities $\boldsymbol{v}^{\text{lin,ang*}}_{t}$ and stance signal $\phi_{t}^{\text{stance}}$, where t=1,2,,Tt=1,2,...,T and $T$ is the sequence length:
(ii) Upper-Body Joints Tracking Error: $E_{\text{tracking}}^{\text{upper}}(\boldsymbol{q}^{\text{upper*}}_{t})=\frac{1}{T}\sum_{t=1}^{T}\left|\boldsymbol{q}_{t}^{\text{upper}}-\boldsymbol{q}_{t}^{\text{upper*}}\right|$
(iii) Root Velocity Tracking Error: $E_{\text{tracking}}^{\text{root}}(\boldsymbol{v}_{t}^{\text{lin,ang*}})=\frac{1}{T}\sum_{t=1}^{T}\left|\boldsymbol{v}_{t}^{\text{lin,ang}}-\boldsymbol{v}_{t}^{\text{lin,ang*}}\right|$
@@ 4.2 Baselines
    Upper-PD-w/o-Force-Curr.: A baseline following [lu2024mobile, homie], using PD control for upper-body joint tracking without force randomization.
@@
    Upper-PD-ID: Extends (a) with a learned force estimator [portela2024learning] and inverse-dynamics-based torque compensation under quasi-static assumptions (details in Section A.3).
@@
    Monolithic-WB-RL-w/o-Force-Curr.: Built upon prior designs [cheng2024expressive, he2024hover], a single agent is trained with the same goal commands as FALCON, but without applying any force during training.
@@ 4.3 Simulation Results
We evaluate FALCON and baselines on 252 ACCAD [accad] motion targets under three force levels: (i) No-Force ($\alpha_{g}=0$), (ii) Middle-Force ($\alpha_{g}=0.5$), and (iii) Large-Force ($\alpha_{g}=1.0$), applied to both end-effectors. As shown in Table 1, across all settings, FALCON with force curriculum achieves the lowest tracking errors in both upper-body motion ($E^{\text{upper}}_{\text{tracking}}$) and root velocity ($E^{\text{root}}_{\text{tracking}}$), demonstrating robust manipulation under disturbance. Under L-Force, it reduces upper-body error to 0.37, outperforming PID-Force-Curr. (0.60) and M-WB-RL (0.73). Root error remains low at 0.45, indicating stable locomotion. While force curriculum benefits all methods, FALCON gains most due to its decomposed learning structure.
@@ 4.4 Real-World Quantatitive Tracking Results
We evaluate FALCON on Unitree G1 with each hand loaded with 1.2kg payload in a real-world task, which is walking at (0.5,0.0)(0.5,0.0)m/s with zero angular velocity, fixed height and waist, and keeping the upper body in its default position. We compare against two baselines: (i) Upper-PD with Force Curriculum, and (ii) Monolithic-WB-RL with Force Curriculum. As shown in Table 4, FALCON achieves the lowest tracking errors, and perform stable and natural motion in heavy-duty loco-manipulation.